### v2

# The trichotomy of primordial black holes initial conditions

Abstract
We show that the threshold to form a black hole, in an asymptotically flat and radiation-dominated Friedman-Robertson-Walker (FRW) Universe, is not solely (mainly) determined by the behaviour of the compaction function at its extrema, as earlier thought, but also by the Ricci scalar of the spatial geometry at smaller (but super-horizon) scales, which we call “the core”. We introduce three classes of initial conditions characterised by an open (O), closed (C), or flat (F) FRW core surrounded by a shell with higher three-dimensional curvature. In the C case, the core helps the collapse so that the black hole formation threshold is the lowest among all cases. Type-II black holes might only be generated by Type-O or F (each of those with different thresholds, with O being the highest) or by a Type-C with a negligible Ricci scalar at the centre, which we call an effective F core. Finally, we argue that an F core is typically more probable for a sharp power spectrum, however, it is also more likely related to non-spherical initial conditions. On the other hand, a very broad power spectrum, which might be related to the observed NanoGrav signal, would favour the formation of Type-I black holes with a mass spectrum peaked at the Infra-Red scale.
I Introduction
The knowledge of the threshold leading to black hole formation is a key ingredient in predicting their abundance in our Universe. In a seminal paper ss , Shibata and Sasaki argued that the conditions to form a black hole from a large over-density is fully determined by the behaviour of the compaction function, which is basically the gravitational potential within a sphere of radius $r$, once the infinitely long wavelength part of the perturbation is extracted. In the case of Type-I black holes, where the compaction function has only one maximum, Musco musco set the threshold to trigger gravitational collapse to the magnitude of this value. Although this threshold does depend upon the full shape of the compaction function musco , it has been argued in universal that it is mainly determined by its curvature at the maximum. With this at hand, an analytical formula, matching up to a few percent error, the numerical results albert , has been provided. This opened up the doors to construct the non-linear statistics of primordial, Type-I, black holes sheth .
Type-II black holes instead, are characterised by the presence of a minimum in the compaction function, surrounded by two maxima. The distinction between this and the Type-I case is mathematically determined by the magnitude at the maximum ($r_{m}$) of a variable constructed from the linear smoothed over-density multiplied by the square of the background areal radius. This variable, dubbed as $g$, gained a central role in the statistics of Primordial Black Holes (PBHs) abundances sheth . Statistically, $g$ inherits the probability distribution of curvature perturbations.
Until very recently, Type-II black holes were discarded as they needed larger values of $g$ ($\frac{4}{3}$) than those of Type-I, and this was thought to be unlikely. Nevertheless, in a recent paper jacopo , the Authors showed that, in the case of a broad power spectrum, the non-linear statistics of Type-I profiles, developed by the use of the analytical formula of universal , prefer threshold values extremely close to Type-II amplitudes.
Shortly after the breakthrough on numerical simulation of Type-II black holes T2 , which resolved the coordinate singularity found in kopp , Escrivà in w , showed that for certain curvature profiles of the linear smoothed over-density, parametrised by their dimensionless curvature at $r_{m}$ ($w$), would produce Type-I or II black holes dependently upon the size of $w$. This shed doubts on the validity of the analytical formula universal , which, for Type-I, was supposed to be valid for any value of $w$.
What we are going to argue here is that there is a trichotomy of possible initial conditions that provides three different behaviours for the PBH formation thresholds:
Consider an isolated spherically symmetric scalar perturbation. At super-horizon scales, a local patch of the perturbed Universe can be approximated by a Friedman-Robertson-Walker (FRW) metric with certain constant three-dimensional curvature (separate Universe). This patch, which we call the core, is either a closed (Type-C), an open (Type-O) or a flat (Type-F) FRW Universe. By looking at larger distances, however, the spacetime becomes more and more inhomogeneous until reaching a peak of the co-moving three-dimensional curvature
${\cal R}\equiv a^{2}\times{}^{(3)}R\$, (1)
which in turn is related to a shell with the highest non-linear over-density. This over-dense shell is the one that, if over-threshold, eventually collapses forming a black hole.
What we empirically found is that, while Type-O and F would not follow the analytical formula of universal for large values of $w$, as already noted in w for the F case, while a Type-C fluctuation, with a non-negligible core, would instead follow it increasingly better for higher values of $w$.
In other words, the three core cases provide three distinct thresholds for the maximum of $g$, with the C type threshold being the lowest. The reason of it is that in the C case, the core geometry would help the collapse. Whether this extra ingredient is relevant or not depends on the spatial extension and amplitude of the internal core, which is related a) to the sharpness of the surrounding shell parametrised by $w$ and b) by its amplitude:
[…]
II Types of initial conditions
Working beyond the separate universe approach, within gradient expansion, a spherically symmetric metric at super horizon scales111i.e. for wavelengths larger than $(aH)^{-1}$, where $a$ is the scale factor of the background FRW Universe and $H\equiv\frac{\dot{a}}{a}$. can be written as
$ds^{2}=-dt^{2}+a^{2}(t)e^{2\zeta(r)}\left(dr^{2}+r^{2}d\Omega^{2}\right),$ (2)
or, equivalently
$ds^{2}=-dt^{2}+a^{2}(t)\left(\frac{dx^{2}}{1-K(x)x^{2}}+x^{2}d\Omega^{2}\right).$ (3)
Using the first set of coordinates, we can write the compaction function ss , $\mathcal{C}(r)$, in terms of the curvature fluctuation yoo (in units of $G_{N}=1$)
$\mathcal{C}(r)=\frac{2}{3}\left[1-\left(1+r\zeta^{\prime}(r)\right)^{2}\right]=g(r)\left(1-\frac{3}{8}g(r)\right),$ (4)
where we have defined, as in sheth , $g(r)=-\frac{4}{3}r\zeta^{\prime}(r)$, $r=0$ is the locus of the over-density peak (for a maximum) or trough (for a minimum). Mathematically, this variable can also be written in terms of the smoothed linear over density up to a radius $r$ ($\delta_{r}$) sheth , i.e. $g\equiv a_{i}H_{i}^{2}r^{2}\delta_{r}$, where $H_{i}$ and $a_{i}$ (set to be $1$), are the initial Hubble and scale factors. The compaction function represents the gravitational potential enclosed in a radius $r$ generated by the non-linear over-density when the background energy-density is extracted.
The maximum of $g$, which we shall name $r_{m}$, corresponds to a maximum of the compaction function for $g(r_{m})<\frac{4}{3}$ (Type-I black holes), and a minimum for $g(r_{m})>\frac{4}{3}$ (Type-II). For Type-I black holes, the threshold is typically given in terms of the maximum of the compaction function, as there, the gravitational potential is the highest. In Type-II black holes, however, the minimum of the compaction function is surrounded by two maxima. The compaction function within the two maximums, being under-critical, fights against collapse, thus, the formation of a black hole is triggered whenever the magnitude of the non-linear over-density at $r=r_{m}$ is smaller than a threshold. In terms of the linear smoothed over-density at $r_{m}$, however, both cases would require $g(r_{m})>g_{c}$, where $g_{c}$ is profile dependent musco .
Nevertheless, for Type-I black holes, it has been found in universal that $g_{c}$ mostly depends on the physics around the collapsing shell surrounding the largest over-density value (the shell from now on). Specifically, the threshold has been found to depend upon the curvature of the smoothed linear over-density at its maximum, i.e., it would only depend on
$w\equiv-r^{2}\frac{d^{2}g(r)}{dr^{2}}\Big|_{r=r_{m}}\ ,$ (5)
which is its dimensionless curvature at $r_{m}$.
On the contrary, and this is what we are going to argue, in Type-II configurations, the central value of the over-density plays a non-trivial role for the threshold.
II.1 The core
According to gradient expansion, at super-horizon scales, a local patch of the spacetime is well approximated by a FRW Universe with a specific three-dimensional curvature, and the initial conditions for PBH formation can be classified by the sign of ${\cal R}(x\rightarrow 0)\sim 6K(x\rightarrow 0)$, as in the following table:
${\cal R}(x\rightarrow 0)$ Type “small scales” FRW Universe
+ C Closed
0 F Flat
- O Open
Primordial black holes are then formed by the next-to-leading order inhomogeneities in the three-dimensional curvature.
Because we shall mainly work with the metric (2), we need the following relation between $x$ and $r$:
$x=re^{\zeta(r)}.$ (6)
For Type-II profiles, (6) cannot be inverted as it passes through an extrema. Thus, we consider instead
${\cal R}(r)=-\frac{2e^{-2\zeta(r)}\left(2r\zeta^{\prime\prime}(r)+r\zeta^{\prime}(r)^{2}+4\zeta^{\prime}(r)\right)}{r}.$ (7)
[…]
As we can already anticipate, the amplitude and size of the core can either help or resist the collapse of the shell. It indeed turns out, as we are going to see, that Type-II configurations are related to the case in which the core cannot be forgotten in the computation of the threshold.
II.2 Initial conditions
The profiles considered in w , which for large values of $w$ implied Type-II configurations, correspond to Type-F. To make our point, we have taken the exponential (Type-F) profile as a base and then deformed it in order to obtain a closed or open FRW core. In this sense, what we are going to consider are collapsing shells characterised by the three possible configurations for the cores and connect those with Type-I and II black holes.
Starting from the base, we then consider:
  • Type-F:
    $\zeta_{F}(r)=\mu\ e^{-(\frac{r}{r_{m}})^{\beta}}.$ (8)
    The parameter $\beta$ in (8) is related to the dimensionless curvature of $g(r)$ at $r_{m}$ ($w$).
    It is easy to see that in $r\rightarrow 0$, ${\cal R}\rightarrow 0$
    ${\cal R}(r)\approx 4\mu\ \beta(1+\beta)\frac{e^{-2\mu}}{r_{m}^{\beta}}r^{\beta-2}+\mathcal{O}(r^{2\beta-2}),$ (9)
    valid for $\beta\geq 2$. This is a physical condition, as for smaller $\beta$ the Poisson equation would diverge at the centre.
  • Type-C:
    $\zeta_{C}(r)=\mu\left(1-\frac{\left(\lambda\frac{r}{r_{m}}\right)^{2}}{\alpha\left[1+\left(\lambda\frac{r}{r_{m}}\right)^{3}\right]}\right)e^{-(\lambda\frac{r}{r_{m}})^{\beta}},$ (10)
    where $\lambda$ is the solution of
    $\left.\frac{dg(r)}{dr}\right|_{r=r_{m}}\propto\left.\zeta^{\prime}(r)+r\zeta^{\prime\prime}(r)\right|_{r=r_{m}}=0,$ (11)
    enforcing that $r_{m}$ is the maximum of $g(r)$.
    The expansion of (7) for this profile corresponds to a closed FRW Universe:
    ${\cal R}(r)\approx 24\mu\frac{e^{-2\mu}\lambda^{2}}{\alpha r_{m}^{2}}+\mathcal{O}(r^{\beta-2}),$ (12)
    valid for $\beta\geq 2$.
    As we can see from Eq.(12), the role of the $\alpha$ parameter is to suppress the core amplitude of the perturbations.
  • Type-O:
    $\zeta_{O}(r)=\mu\left(1+\frac{\left(\lambda\frac{r}{r_{m}}\right)^{2}}{\alpha\left[1+\left(\lambda\frac{r}{r_{m}}\right)^{3}\right]}\right)e^{-\left(\lambda\frac{r}{r_{m}}\right)^{\beta}}.$ (13)
[…]
    ${\cal R}(r)\approx-24\mu\frac{e^{-2\mu}\lambda^{2}}{\alpha r_{m}^{2}}+\mathcal{O}(r^{\beta-2}),$ (14)
    is valid for $\beta\geq 2$.
[…]
Figure 1: Typical ${\cal R}(r)$ for the three types of perturbation at threshold. All of them correspond to $w\simeq 38.6$. The dashed vertical line is $r=r_{m}$.
[…]
III Numerics
To check whether an initial perturbation collapses or not, and thus find the threshold in terms of $g(r_{m})$, we have used the public SPriBHoS-II code codeII . The threshold would initially be given in terms of a critical $\mu=\mu_{c}$ and then translated in terms of $g(r_{m})$.
To track the black hole formation more efficiently, we have modified the code, including an auxiliary function defined as twice the Misner-Sharp mass $M$ ms over the areal radius $R$
$\mathcal{C}_{0}=2\frac{M}{R}\ .$ (15)
An apparent horizon leading to a black hole is only formed when $\mathcal{C}_{0}$, at its maximum, reaches $1$ twice. The first $1$ would signal the cosmological horizon crossing of the super-horizon perturbation. After this, because of the dilution of the Universe expansion, $\mathcal{C}_{0}$ would decrease. However, over threshold, the decay would then be followed by a growth, and a sub-horizon trapped surface would form whenever $\mathcal{C}_{0}\sim 1$ for the second time.
Large values of $w$ (>30>30), i.e. initial sharp profiles, required a substantial modification of the number of Chebyshev points ($N_{\text{cheb}}$) in the pseudo-spectral method used in w . The code becomes increasingly unstable for $w>50$ in the Type-C case. However, as we shall show, the threshold already saturates at those values with no need of extra probes [^2] In fact, Type-C configurations for $w>50$ when simulated with the SPriBHoS-II code, show a clear trend to the formation of a trapped horizon with a saturated threshold. This happens way before the simulation breaks because of Gibbs phenomena..
Finally, simulations are done in units $G_{N}=1$ with an initial Hubble scale $H_{i}=\frac{1}{2}$. All our numerical results will then be given in those units.
See Appendix B for the evolution of the Hamiltonian constraint.
IV Thresholds
Here we present the threshold values, $g_{c}$, for different profiles. In Fig.2, we can clearly see that as $w$ increases, the trichotomy appears. For Type-F and Type-C perturbations, the crossing between Type-I and Type-II happens around $w\sim 34$, while for Type-O happens before, at $w\sim 29$. We also see the trend anticipated earlier, that the lowest thresholds are related to Type-C, where the cores help the collapse. On the other hand, the highest thresholds are related to Type-O, where instead the cores fight the collapse.
This is one of the main results of our paper, namely that for large enough cores, the threshold cannot only depend upon the shell structure.
Figure 2: Numerical threshold, in terms of $g$, for the three types of perturbations. The dashed line corresponds to $g=\frac{4}{3}.$ All the points correspond to different configurations of the $\alpha$ and $\beta$ parameters (see Appendix A).
For large values of $w$ (still in Fig. 2), Type-O and Type-F become Type-II, as already noted in w . On the other hand, Type-C perturbations lie in both Type-I and Type-II regions. We will explore this peculiar behaviour in the following sections, nevertheless, we can already anticipate that Type-C should smoothly behave as Type-F for small core amplitudes. This is precisely what happens in the Fig. 2.
IV.1 Type-I
The threshold for Type-I perturbations suggested by universal was computed in terms of the dimensionless curvature of the compaction function ($q$)
$q=\frac{-r_{m}^{2}\partial_{r}^{2}\mathcal{C}(r_{m})}{4\mathcal{C}(r_{m})\left(1-\frac{3}{2}\mathcal{C}(r_{m})\right)}.$ (16)
In terms of $\cal C$ at its maximum, it was then found to be
$\mathcal{C}_{c}(q)=\frac{4}{15}e^{-\frac{1}{q}}\frac{q^{1-\frac{5}{2q}}}{\Gamma(\frac{5}{2q})-\Gamma(\frac{5}{2q},\frac{1}{q})}.$ (17)
which matches within a few percent error the threshold found in numerical simulations.
As we already mentioned, we shall work instead with $g$, which is more convenient for Type-II perturbations. Notice, however, that because the relation between $\mathcal{C}$ and $g$ is quadratic (4), the error in terms of $g$ is generically higher than that of $\mathcal{C}$ (see Fig.4).
In Fig. 3, we selected, from Fig.2, all Type-I configurations.
Figure 3: Numerical threshold values, in terms of $g$, for Type-I perturbations.
Figure 4: Relative deviation, $d$, for Type-I perturbations, considering the analytical formula (17). The left panel is the relative deviation in terms of $g$ and the right panel in terms of $\mathcal{C}$. Here, one can clearly see that for all profiles considered, the relative deviation $d_{\mathcal{C}}<2\%$.
There, the relative deviation between the analytical threshold ($\delta_{A}$) and the numerical one ($\delta_{N}$), $d\equiv|\delta_{N}-\delta_{A}|/\delta_{N}$, is $\sim 5\%$ (which reduces to $2\%$ for $\mathcal{C}_{c}$) for values of $w\sim 16$ and decreases to $\sim 1.8\%$, as $w$ increases ($0.15\%$ for $\mathcal{C}_{c}$), c.f. Fig.4.
IV.2 Type-II
Type-II perturbations require a higher threshold with respect to those of Type-I, as there the core would not help for the collapse of the external (higher density) shell. Illustrative examples are given in Fig.1, where Type-II is obtained by lowering the core from a Type-C(I).
In w it was shown that, at order $w\sim 30$ all the initial profiles considered (that can be checked to be of Type-F), at threshold, would produce Type-II black holes. In Fig.3, we have explicitly shown that this is not always the case for Type-C, while it is true for Type-F and O. Nevertheless, there is also a distinction between those two. Considering the fitting formula proposed in w , i.e.
$g_{c}=aw^{b}\hskip 28.45274pt(a\approx 0.51962,\ b\approx 0.26687),$ (18)
for $w\gtrsim 30$, one can see that it would only work for Type-F. Testing (18) for flat profiles (8) leads to a relative deviation with respect to the numeric of less than $0.3\%$. However, for Type-O, this analytical formula does not fit so well the numerical values, leading to a relative deviation of $\sim 5\%$. For those Type-C profiles that cross to the Type-II region, the proposed analytical formula gives us a full range of values for similar values of $w$, from $d\sim 10\%$, which clearly largely deviate from the analytical formula, to $d\sim 2\%$ (see Fig.5).
Figure 5: The left panel is the numerical threshold, in terms of $g$, for Type-II perturbations. The right panel is the relative deviation, $d$, in terms of $g$ for Type-II perturbations using (18).
[…]
V Transition between closed and flat perturbations
As we have shown, Type-C perturbations are characterised by constant strictly positive core values for ${\cal R}(x)$ or ${\cal R}(r)$ (Fig.1).
One may ask what happens if this constant value decreases approaching zero, if it would still behave as Type-C or rather as a Type-F [^3] We thank Jaume Garriga for posing this question..
As we have argued, the trichotomy appears as the shell gets thinner, and the core plays a role. Then, if the core corresponds to a closed FRW, it helps the collapse of the perturbation. In this case, a “low” threshold is in order, leading to a Type-I BH. On the other hand, in the flat case, the core does not help the collapse, implying a higher threshold leading to Type-II BHs.
Because of the central importance of the size of the collapsing shell, we propose a new dimensionless parameter:
$w_{\cal R}\equiv-r^{2}\frac{\partial^{2}_{r}{\cal R}(r)}{{\cal R}(r)}\Big|_{r=r_{p}}\ .$ (19)
In Fig.6, we can see that, fixing $w_{\cal R}$ we can go to a Type-II by lowering the core amplitude from a Type-I configuration with closed core (which we call Type-C(I)). Once the perturbation becomes Type-II, which requires a higher threshold, the core effect is negligible, and hence, it effectively acts as a flat one, we call this Type-C(II) configuration.
[…]
As we are going to argue, the core plays an important role if it fulfils the following two conditions for ${\cal R}(r)$:
  • The amplitude of the core, ${\cal R}(0)$, is large enough with respect to the peak amplitude, ${\cal R}(r_{p})$,.
  • The peak is sharp enough, meaning $w_{\cal R}$ is large enough.
The first condition is actually connected to the second one and can be tested by computing the ratio
${\cal Q}\equiv\frac{{\cal R}(0)}{{\cal R}(r_{p})}\Big|_{c},$ (20)
[…]
Figure 7: Values for ${\cal Q}\equiv\frac{{\cal R}(0)}{{\cal R}(r_{p})}\Big|_{c}$ for different values of $w_{\cal R}$. The blue points correspond to Type-I perturbations, while the red ones to Type-II. The line corresponds to a tentative transition region (Eq.(21)) that depends on the value of $w_{\cal R}$.All the data points correspond to the different profiles shown in Appendix A.
If the peak is broad enough (i.e. $w_{\cal R}$ small), we can consider it as a closed FRW, which will help the collapse. Then, the core behaviour is not noticeable. That is the reason why, for small values of $w_{\cal R}$ the three types of perturbations are Type-I. On the other hand, for higher values of $w_{\cal R}$, the peak gets thinner and loses relevance with respect to the core.
The transition region between Type-I and II black holes obtained in Fig.7 follows the tentative formula:
${\cal Q}=0.70-\frac{70.78}{(w_{\cal R})^{1.48}}.$ (21)
Once we have identified the region of Type-C perturbations with a negligible core, we can interpret the Type-C(II) points as effective Type-F ones (those below the curve $\cal Q$, c.f. Fig.8).
Figure 8: Numerical threshold, in terms of $g$ as a function of $w$, for the three types of perturbations, including the Type-C(II) profiles that behave effectively as flat. The dashed line corresponds to $g=\frac{4}{3}$. All the data points correspond to the different profiles shown in Appendix A.
For $w_{\cal R}\lesssim 22$ ($w\lesssim 30$), where ${\cal Q}=0$, all configurations lead to Type-I black holes while, for $w_{\cal R}\xrightarrow$$\infty$, our fitting formula for $\cal Q$ indicates that, in order to have Type-I configuration one needs a core amplitude to be at least $\sim 70\%$ of the peak one.
[…]
Figure 9: Numerical threshold, in terms of $g$ as a function of $w_{\cal R}$, for the three types of perturbations, including the effective flat profiles. The dashed line corresponds to $g=\frac{4}{3}.$
VI Initial conditions statistics
PBHs are formed by rarely large smoothed over-densities on a radius $r$ sheth . The threshold is then typically given in terms of the compaction function, which, if the perturbation is spherical, is the combination (4) of the statistical variable sheth
$g(r,\vec{x}_{0})=-\frac{4}{9}\int\frac{d^{3}k}{(2\pi)^{3}}e^{i\vec{k}\cdot\vec{x_{0}}}(kr)^{2}W(kr)\zeta_{k},$ (22)
where $r=|\vec{x}-\vec{x}_{0}|$ and $\zeta_{k}$, assumed to be independent and Gaussian-distributed, are the Fourier modes of the curvature perturbations in synchronous gauge and at leading order in gradient expansion. Note that here we have not yet assumed any spherical symmetry and
$W(y)=3\frac{\sin(y)-y\cos(y)}{y^{3}}\ ,$ (23)
is the Fourier transformation of the Heaviside theta function. For spherically symmetric configurations of the over-density, as we already pointed out, $g(r,\vec{x}_{0})=a_{i}H_{i}^{2}r^{2}\delta_{r}$, and the combination (4) indeed corresponds to the effectively asymptotically flat “gravitational potential” of the perturbation.
Typically, one expects that a black hole is formed by a large over-density (peak) at the point $\vec{x}=\vec{x}_{0}$, which statistically implies an almost spherical symmetry around $\vec{x}_{0}$ bbks . In this case, without any loss of generality, one can fix $\vec{x}_{0}$ as the origin of the spherical symmetry.
Rewriting (7) in terms of $g$ and expanding it near the origin one finds
$\frac{{\cal R}(r)}{H_{i}^{2}}=\left.\frac{3}{8}\frac{e^{-2\zeta(r)}\left(8g(r)-3g(r)^{2}+8rg^{\prime}(r)\right)}{r^{2}{H_{i}^{2}}}\right|_{r\xrightarrow{}0}\approx 3e^{-2\mu}\lim_{r\rightarrow 0}\frac{g(r)}{r^{2}H_{i}^{2}}=3e^{-2\mu}\lim_{r\rightarrow 0}\delta_{r}\simeq\delta_{0}\ ,$ (24)
where $\delta_{0}$ is the central over-density [^4] Note that, at the origin, the smoothed linear over-density is proportional to the non-linear over-density. and we have used the empirical fact that $3e^{-2\mu}\sim\mathcal{O}(1)$ at threshold.
We then already see that in the case of Type-F configuration, where $\delta_{0}=0$, instead of a peak of the over-density we are in the presence of a trough (a minimum). Thus, statistically, Type-C configurations (or effectively Type-F) are hardly spherical michiro . In this case, the threshold for PBH formation will be larger than the one discussed above, implying an extra statistical suppression for their realisation with respect to the Type-C case. On the other hand, as we are going to see, the Type-C(II) core is less statistically suppressed with respect to the Type-C(I) one.
The estimation of which core type is finally most relevant for the PBHs abundance would then require a better statistical and numerical understanding of non-spherically symmetric configurations, although the numerical evolution of those is already an active field of research nonsph . Nevertheless, in certain limiting configurations of the power spectrum, the selection of the specific initial conditions might be predicted without this knowledge, as we shall see.
The variable $\delta_{0}$  [^5] Earlier statistics on PBHs abundances were constructed on this variable, e.g. mg ., as $g$, is multi-Gaussian with variance jacopo ,
$\sigma_{\delta_{0}}^{2}=\frac{4}{81H_{i}^{4}}\int dk\ k^{3}\mathcal{P}(k)\ ,$ (25)
where $\mathcal{P}(k)$ is the dimensionless power spectrum of the curvature perturbations $\zeta_{k}$.
For illustrative purposes, we will consider the two typical examples of a very narrow and a very broad $\mathcal{P}(k)$:
  • Narrow distribution, $\mathcal{P}(k)=\mathcal{A}_{s}k_{p}\delta(k-k_{p})$.
    $\sigma_{\delta_{0}}^{2}=\frac{4}{81H_{i}^{4}}\mathcal{A}_{s}k_{p}^{4}.$ (26)
  • Broad distribution, $\mathcal{P}(k)=\mathcal{A}_{s}\theta(k-k_{IR})\theta(k_{UV}-k)$, with $k_{UV}\gg k_{IR}$.
    $\sigma_{\delta_{0}}^{2}=\frac{1}{81H_{i}^{4}}\mathcal{A}_{s}k_{UV}^{4}.$ (27)
In Fig.10 we see that at threshold, the core values for Type-C(I) are of $\sim\mathcal{O}(10^{-5}-10^{-6})$ for all the cases studied.
Figure 10: Values of ${\cal R}(r)$ at the centre for Type-C perturbations.
Let us now consider that the peak of the PBH mass spectrum is associated to a specific Fourier mode $k_{\bullet}$. The order of magnitude of the PBH mass ($M_{\bullet}$) is musco $M_{\bullet}\sim\frac{M_{p}^{2}}{H_{k}}$, where $H_{k}$ is the Hubble scale when the mode $k$ crossed the horizon. One then finds
$k_{\bullet}\sim M_{p}\sqrt{\frac{H_{i}}{M_{\bullet}}}\ .$ (28)
We start by assuming that $k_{UV}$ and $k_{p}$ are of order $k_{\bullet}$, which is what is expected for a very sharp power spectrum jacopo and (optimistically jacopo ), $\mathcal{A}_{s}\sim 10^{-1}$, then
$\sigma^{2}_{\delta_{0}}\sim 10^{-2}\frac{M_{p}^{2}}{M_{\bullet}^{2}}\ .$ (29)
To maximise $\sigma^{2}_{\delta_{0}}$, we consider the lowest asteroid mass PBH (unconstrained to be the whole of dark matter Carr ). In this case $M_{\bullet}\sim 10^{41}\text{ GeV}$ and so
$\sigma^{2}_{\delta_{0}}\sim 10^{-46}.$ (30)
Then,
$\sqrt{\sigma_{\delta_{0}}^{2}}\ll\delta_{0}(\text{Type-C(I)})\sim 10^{-5},$ (31)
implying a statistical suppression of Type-C(I) profiles with respect to Type-F, since one should go almost $10^{18}$ variances away from the central value to obtain Type-C(I) perturbations.
For a very broad power spectrum, however, the peak of the mass distribution might instead be related to the IR scale, as it would be in the presence of only Type-I black holes jacopo . In that case, for Type-C(I) not to be statistically suppressed, one would need $k_{UV}/k_{IR}>10^{9}$ to obtain $\sqrt{\sigma_{\delta_{0}}^{2}}\sim\delta_{0}(\text{Type-C(I)})$. Then, the analysis of jacopo would follow since Type-C(I), given their lower threshold, would dominate the abundance of PBHs.
We note that in NANOGrav , such a large separation of scales were considered when interpreting the NANOGrav signal nano as a second-order Gravitational Waves background sourced by large curvature perturbations.
VII Conclusions
According to the separate Universe approach, a local super-horizon patch (the core) of a perturbed Universe is described by a FRW geometry with a certain three-dimensional curvature. At larger scales, instead when the inhomogeneities build up, the typical configuration related to a PBH formation is a shell with an over-density peak. The formation of a BH happens whenever the amplitude of this shell is high enough for the gravitational collapse to fight against the Universe dilution. If the core is a closed Universe (Type-C) configuration, it helps the collapse of the shell, otherwise, either it fights against it (Type-O(pen)), or it is irrelevant (Type-F(lat)). Therefore, for Type-F and O configurations, the required amplitude of the energy density to form a black hole (the threshold) is bound to be higher than that of Type-C. In this sense, there is a trichotomy of possible initial conditions, each one having a different threshold for BH formation depending upon the three-curvature peak and core amplitudes. In turn, this shows that the relevant variable for understanding BH formation in an expanding Universe is rather the non-linear over-density instead of the compaction function or the linear over-density, as thought earlier.
This trichotomy is, however, only relevant if the core is extended enough to interfere with the shell. This empirically happens if the dimensionless curvature of the over-dense shell $w_{\cal R}\equiv-r^{2}\frac{\partial_{r}^{2}{\cal R}}{{\cal R}}$ calculated at the peak $r=r_{p}$, is higher than $\sim{\cal O}(17)$ ($\sim{\cal O}(30)$ for $w$), which is when the first point of Type-O crosses the Type-II region. For broader cases, the type of black hole formed is always of Type-I. However, when the core becomes relevant, only Type-C configuration with large enough core amplitudes may generate PBHs of Type-I. In all other cases, Type-II black holes will always be formed. For Type-C, the separatrix between Type-I and Type-II core amplitudes has been empirically found in eq. (21), and depends upon the ratio of the core amplitude with respect to the peak one, being parametrised by $w_{\cal R}$. For very sharp shells, we found that the core amplitude must be at least about $70\%$ the peak one in order to form Type-I black holes.
A more extensive and complete analysis should be done considering different perturbation profiles, $\zeta(r)$, which we leave for future work.
Finally, we argued that, at least for Gaussian distributed curvature perturbations, shapes with zero curvature cores (Type-F in spherical configuration) are favourable. However, whether black holes will be more likely related to these shapes is a different question, which can only be answered with the statistics of thresholds. For example, Type-F configurations are also likely to generate a highly non-spherical initial condition for the non-linear over-density (note that this would be statistically different from the linear one), making gravitational collapse more difficult than in the spherical case. The reason is that a sharp shell can be thought of as generated by the interference of an infinite number of peaks, which would distort the spherical symmetry. For highly non-spherical configurations, a large part of the energy density of the over-dense shell would be dissipated away in the form of gravitational waves, thus a larger threshold.
For a very broad power spectrum, however, which might be related to the NanoGrav gravitational waves observations NANOGrav , the statistical preference of Type-F with respect to Type-C(I) becomes negligible, in that case $\sqrt{\sigma_{\delta_{0}}^{2}}\sim\delta_{0}(\text{Type-C(I)})$. In this case, the higher threshold necessary to form Type-F black holes would favour Type-C(I) for the PBHs abundance. This, in turn, would imply a black hole’s mass spectrum peaked at the Infra-Red scale of the power spectrum jacopo , instead of at the Ultra-Violet scales, as typically assumed.
Whether for a generic power spectrum, the abundance of PBHs is finally dominated by Type-I black holes, as suggested in jacopo , or not, is however still an open question which, to be answered, requires further (physical and statistical) knowledge of the non-spherical gravitational collapse. This interesting problem is left for future studies.
Acknowledgements.
The authors are grateful to Albert Escrivà, Jacopo Fumagalli, Jaume Garriga, Cristian Joana and Ravi Sheth for discussions on thresholds, profiles and Types-I/II initial conditions. The research of CG is supported by the grant PID2022-136224NB-C22, funded by MCIN/AEI/10.13039/501100011033/FEDER, UE, by the grant/ 2021-SGR00872 and CEX2024-001451-M funded by MICIU/AEI/10.13039/501100011033.
Appendix A Profile parameters
Here, we provide all the parameters used for each profile shown.
$\mu$ $w$ $w_{\cal R}$ $\lambda$ $\alpha$ $\beta$ $\mu$ $w$ $w_{\cal R}$ $\lambda$ $\alpha$ $\beta$
0.678 16.11 6.45 0.939 1 5.53 0.605 38.69 30.68 0.983 1.74 6
0.671 23.09 17.33 0.953 1 6.075 0.616 39.32 32.36 0.981 1.5 6.2
0.613 30.11 22.79 0.981 1.8 5.5 0.602 39.34 31.04 0.984 1.8 6
0.600 33.16 25.33 0.984 2 5.7 0.658 40.60 40.15 0.969 1 7.1
0.665 33.16 31.05 0.964 1 6.7 0.634 40.60 36.76 0.974 1.1542 6.7
0.551 33.81 23.42 0.998 13 5.1 0.636 42.10 37.74 0.976 1.2 6.7
0.581 34.08 24.88 0.990 3 5.4 0.610 42.32 34.28 0.983 1.6 6.2
0.571 34.12 24.44 0.993 4 5.3 0.634 42.71 38.10 0.976 1.2175 6.7
0.558 34.45 24.16 0.996 7 5.2 0.658 42.71 42.51 0.970 1 7.2
0.662 34.87 33.27 0.965 1 6.8 0.619 43.56 36.93 0.980 1.4 6.5
0.565 34.88 24.83 0.994 4.7 5.3 0.645 44.02 41.41 0.974 1.1 6.99
0.600 35.30 27.02 0.985 2 5.7 0.635 44.23 39.83 0.977 1.2 6.8
0.633 35.21 29.73 0.975 1.3 6.2 0.657 44.60 44.62 0.971 1 7.29
0.578 35.92 26.34 0.990 3 5.5 0.629 45.95 40.77 0.978 1.25 6.8
0.586 36.15 26.98 0.988 2.5 5.6 0.638 46.31 42.74 0.976 1.15 6.99
0.661 36.74 35.56 0.966 1 6.9 0.644 46.43 43.85 0.975 1.1 7.1
0.618 37.36 30.54 0.980 1.5 6.1 0.634 46.67 42.15 0.978 1.2 6.91
0.590 38.52 29.07 0.987 2.2 5.8 0.643 48.67 46.08 0.976 1.1 7.2
0.659 38.62 37.84 0.968 1 7
Table 1: Parameters used for Type-C(I) profiles in Fig.3 and Fig.4 and the corresponding $w$ and $w_{\cal R}$.
Type $\mu$ $w$ $w_{\cal R}$ $\beta$
I 0.595 16.01 10.17 3.8
I 0.565 23.10 13.50 4.3666
I 0.540 33.21 16.53 5.005
II 0.531 38.56 26.47 5.29
II 0.522 44.61 30.55 5.585
II 0.517 49.48 33.72 5.8
Table 2: Parameters used for Type-F profiles in Fig.3 and Fig.5 and the corresponding $w$ and $w_{\cal R}$.
Type $\mu$ $w$ $w_{\cal R}$ $\lambda$ $\alpha$ $\beta$
I 0.595 16.22 17.66 1.001 70 3.8
I 0.558 23.05 15.37 1.004 10 4.27
I 0.457 28.25 17.73 1.026 10 4.1
II 0.444 33.29 19.17 1.022 1 4.4
II 0.433 38.60 23.24 1.019 4.685
II 0.422 44.57 26.45 1.017 4.98
II 0.417 49.54 29.00 1.016 5.194
Table 3: Parameters used for Type-O profiles in Fig.3 and Fig.5 and the corresponding $w$ and $w_{\cal R}$.
$\mu$ $w$ $w_{\cal R}$ $\lambda$ $\alpha$ $\beta$ $\mu$ $w$ $w_{\cal R}$ $\lambda$ $\alpha$ $\beta$
0.550 34.02 23.51 0.998 15 5.1 0.611 45.63 37.97 0.983 1.5 6.5
0.556 34.74 24.29 0.996 7.8 5.2 0.566 46.21 34.16 0.992 3 6
0.563 35.25 25.10 0.994 5.1 5.3 0.628 46.29 40.95 0.979 1.26 6.8
0.568 35.96 25.85 0.993 4 5.4 0.624 47.55 41.58 0.980 1.3 6.8
0.572 36.86 26.81 0.992 3.4 5.5 0.616 47.96 40.86 0.982 1.4 6.7
0.571 37.08 26.91 0.992 3.5 5.5 0.633 47.98 43.62 0.978 1.19 6.99
0.576 37.88 27.85 0.991 3 5.6 0.632 48.39 43.83 0.978 1.2 6.99
0.593 41.17 31.97 0.987 2 6 0.564 48.48 35.82 0.993 3 6.1
0.604 42.34 33.97 0.984 1.7 6.2 0.553 48.51 35.00 0.995 4 6
0.598 43.48 34.53 0.985 1.8 6.2 0.663 48.56 42.71 0.978 1.2 6.9
0.617 44.25 37.28 0.981 1.43 6.5 0.632 48.63 44.05 0.978 1.2 7
0.615 44.65 37.49 0.982 1.45 6.5 0.637 48.80 45.14 0.977 1.15 7.1
0.625 45.30 39.54 0.979 1.3 6.7 0.640 49.68 46.61 0.977 1.12 7.2
0.589 45.44 35.42 0.988 2 6.2
Table 4: Parameters used for Type-C(II) profiles in Fig.5 and Fig.9 and the corresponding $w$ and $w_{\cal R}$.
Appendix B Hamiltonian constraint
To test the convergence of the simulations, we have analysed the evolution of the Hamiltonian constraint equation, (as in albert )
$\mathcal{H}=D_{r}M-4\pi\Gamma\rho R^{2},$ (32)
[…]
The results for the evolution of the Hamiltonian constraint using the three profiles considered at threshold (Eq.(8)-(13)), are shown in Fig.11.
Figure 11: Numerical evolution of the Hamiltonian constraint for different profiles at threshold. The time of horizon crossing, when the perturbation re-enters the horizon, is $t_{H}=t_{0}\left(a_{i}r_{m}H_{i}\right)^{2}$, where $t_{0}=1$ for the simulations.
In Fig.11, one can see that around $t/t_{H}\sim 10^{1}$, the Hamiltonian constraint evolution oscillates and diverges from the previous monotonic behaviour. That is because around that point the BH is formed, and then the singularity, hence the numerical evolution does not hold any more.